OpenStudy (anonymous):

***WILL MEDAL*** What is the common difference between successive terms in the sequence? 9, 2.5, –4, –10.5, –17, ... –11.5 –6.5 6.5 11.5

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

the common difference \(a_n=a_1+(n-1)d \\a_2\implies2.5=9+(1)d\\2.5-9=d \\-6.5=d\)

OpenStudy (anonymous):

thank you so much!!